OpenStudy (anonymous):

the longer leg of a 30°-60°-90° triangle is 18. What is the length of the other leg?

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

tn 30 = sqrt3 /3 sqrt3 /3 = AB/18 AB = 18sqrt3 / 3 AB = 6sqrt3